OpenStudy (anonymous):

Could someone help? Medal for best answer. Divide. Assume the denominator does not equal zero.

OpenStudy (anonymous):

\[\frac{ 12x^2y-6xy+4x^2 }{ 2xy }\]

OpenStudy (anonymous):

Please explain how to get the answer, don't just give the answer to me.

OpenStudy (mrnood):

first of all compare it with this: \[\frac{ a+b+c }{ d } = \frac{ a }{ d} + \frac{ b }{ d} + \frac{ c }{ d }\] so break up your expression in th esame way into three terms

OpenStudy (anonymous):

So,\[\frac{ 12x^2y }{ 2xy }+\frac{ -6xy }{ 2xy }+\frac{ 4x^2 }{ 2xy }\]

OpenStudy (mrnood):

good now just cancel factors top and bottom in each term (e.g. cancel out the y in the first term)

OpenStudy (anonymous):

\[6x+(-3)+\frac{ 2x }{ y }\] Like that?

OpenStudy (mrnood):

That seem to be the best option in my opinion

OpenStudy (anonymous):

Is there any more steps to this, or is that the answer?

OpenStudy (mrnood):

well all the terms are of different 'orders' of xy so that is an acceptable answer you could collect the x terms - but this is not really a 'simpler answer: \[\frac{ x(6y+2) }{ y }-3\]

OpenStudy (anonymous):

Thanks for your help :)